# Mission-Aligned Learning-Informed Control of Autonomous Systems: Formulation and Foundations

### v1

@@ 4.2.1 Relational Logic
Syntactically, in first-order logic (FOL), we define a vocabulary (signature) $\Sigma$ consisting of: a set of constant symbols $\mathcal{C}$ (e.g., “Room1”,“Patient2”); a set of variable symbols $\mathcal{V}$ (e.g., position “$x$”, distance“$d$”); a set of function symbols $\mathcal{F}$ (e.g., “$\text{neighborOf}(x)$”), each with an associated arity $n\in\mathbb{N}$ denoted by $f/n$ where $f\in\mathcal{F}$; and a set of predicate symbols $\mathcal{P}$ (e.g., “$\text{Asleep}(x)$” or “$\text{Nursing}(x,y)$”), also with associated arities $P/n$ where $P\in\mathcal{P}$. The set of terms $\mathcal{T}(\Sigma)$ over $\Sigma$ is then defined recursively: $\mathcal{C}\subset\mathcal{T}(\Sigma)$, $\mathcal{V}\subset\mathcal{T}(\Sigma)$; and if $f/n\in\mathcal{F}$ and $t_{1},\dots,t_{n}\in\mathcal{T}(\Sigma)$, then $f(t_{1},\dots,t_{n})\in\mathcal{T}(\Sigma)$ (e.g., “nurseOf(Patient1)”). The set of formulae $\mathcal{F}(\Sigma)$ over $\Sigma$ is then also defined recursively as follows. If $P/n\in\mathcal{P}$ and $t_{1},\dots,t_{n}\in\mathcal{T}(\Sigma)$, then $P(t_{1},\dots,t_{n})\in\mathcal{F}(\Sigma)$ is an atomic formula; if $\phi,\psi\in\mathcal{F}(\Sigma)$, then $\neg\phi,(\phi\land\psi),(\phi\lor\psi),(\phi\rightarrow\psi),(\phi\leftrightarrow\psi)\in\mathcal{F}(\Sigma)$, and if $\phi\in\mathcal{F}(\Sigma)$ and $x\in\mathcal{V}$, then $\forall x\phi$ and $\exists x\phi\in\mathcal{F}(\Sigma)$ (e.g., “$\forall x(\text{Man}(x)\rightarrow\text{Mortal}(x))$”).
@@ 4.2.2 Planning
Conceptually, classical planning could be viewed as a special case of MPC where (i) the transition dynamics ($\mathcal{T}(s,a)$) are discrete and deterministic, (ii) the cost function prioritizes goal achievement (typically 00 if $x\models\mathcal{G}$ otherwise $\infty$), and (iii) the optimization horizon spans until the goal is reached.
@@ 4.4 Planning (and RL) with (Graph and) Relational Neural Networks
Related ideas have also been explored in the reinforcement learning setting with other neuro-symbolic frameworks, such as $\delta$ILP evans2018learning , which aims to relax the problem of Inductive Logic Programming (ILP), or rule learning, into a differentiable setting via (constrained) rule templates where each rule is associated with a learnable weight, similar to the LRNNs, and trained with gradient descent. Building on that, frameworks such as jiang2019neural integrate $\delta$ILP with RL to enhance interpretability and generalization through the logical formalism. The RL problem is modeled as an MDP with logic components $(\mathcal{M},p_{S},p_{A})$, where $\mathcal{M}=(\mathcal{S},\mathcal{A},T,R)$ defines the state space, action space, transitions, and rewards. A state encoder $p_{S}:\mathcal{S}\to 2^{\mathcal{G}}$ maps states to sets of logical atoms, and an action decoder $p_{A}:[0,1]^{|\mathcal{D}|}\to[0,1]^{|\mathcal{A}|}$ translates the logic valuations into action probabilities. Using $\delta$ILP, the policies are represented as valuation vectors $e\in[0,1]^{|\mathcal{G}|}$, where each element indicates the confidence of a logical atom being true. Inference then occurs via a recurrent neural model $f_{\theta}$, which iteratively updates the valuations through time. Logical rules are encoded as weighted first-order clauses of the form
@@ 5.0.1 Fuzzy Sets
where $\mu_{\tilde{A}}(x)=1$ indicates full membership, $\mu_{\tilde{A}}(x)=0$ indicates non-membership, and $10<\mu_{\tilde{A}}(x)<$ 1 represents a partial membership. Common examples of membership functions include triangular, trapezoidal, Gaussian, and sigmoid functions bai2006fundamentals .
@@ 5.0.2 Fuzzy Logic
Fuzzy logic then extends classical Boolean logic by incorporating degrees of truth. In fuzzy logic, the truth value of a statement is again not restricted to binary values (00 or $1$) but can take any value in the interval $[0,1]$. To that aim, fuzzy logic operators extend classical logical operators to operate on the fuzzy sets bai2006fundamentals .
@@ 5.0.3 Fuzzy control
IF $\text{IF }x_{1}\text{ is }\tilde{A}_{1}\text{ AND }x_{2}\text{ is }\tilde{A}_{2}\text{ THEN }y\text{ is }\tilde{B},$ (40)
@@
[…]
@@
IF $\text{IF }x_{1}\text{ is }\tilde{A}_{1}\text{ AND }x_{2}\text{ is }\tilde{A}_{2}\text{ THEN }y=g(x_{1},x_{2}),$ (45)
@@ 6 Integrating the Symbolic and Control Layers
Figure 1: A pictorial model of integrating the Symbolic Planning and Control Layers in Operation. The scheduler decides upon the sequence of Tasks the robot is set to perform. Each Task defines a Planning Problem. The solution to the problem is computed, and then at each time step, for a given state an action is taken which then defines an MPC problem. The MPC problem is solved in real time, and the final (continuous) state is mapped to the set of (discrete) states in the planning ontology. Discrepancies in the form of intermediate rewards and unexpected state transitions can change the planning problem in real-time, necessitating a re-computation of the planning solution.
@@ 6.3 Action-Triggered MPC
$\quad h\left(z_{\tau(k)},v_{\tau(k)};p^{h}(a_{t})\right)\leq 0,\quad v_{\tau(k)}\in\tilde{A}(p^{a}(a_{t}))\,,\quad$ (53c)
@@
In addition, a real-time reward $\hat{R}^{e}(t+1)$ is observed externally at the conclusion of the problem, together with the evaluation of the action closed loop cost functional $\hat{R}^{o}(t+1)$. This quantity $\hat{R}^{e}(t+1)$ represents mission-relevant information that occurred in the course of the operation. For instance, if, in performing one of the control actions, the robot observed negative emotional affect in the human, this can enter into $\hat{R}^{e}(t+1)$ in a manner so as to inform the upper layer, and ultimately the task can be planned to be performed more agreeably to the human, that is, without action $\hat{a}_{t}$.
@@ 6.4 Integrating State representations
where $p_{\text{RobotA}}$ and $p_{\text{LocationB}}$ are their respective positions, and $\sigma$ controls the smoothness of the transition. Note the truncation of the distribution: if the robot is within distance $\bar{d}^{0}_{At}$ of LocationB, we declare that it is At that location with certainty, and if it is a distance at least $\bar{d}^{1}_{At}$ away, then he is crisply declared to not be at the location, with a fuzzy transition function in the annulus defined by the second indicator in the expression.
@@ 6.6 Fuzzy control relaxation alternative
IF $\begin{array}[]{l}\text{IF }\mu_{L_{1}}\text{ is high AND }\mu_{L_{2}}\text{ is low, THEN }a_{t+1}=a_{1},\\ \text{IF }\mu_{L_{2}}\text{ is high, THEN }a_{t+1}=a_{2}.\end{array}$ (62)

### v2

@@ 4.2.1 Relational Logic
Syntactically, in first-order logic (FOL), we define a vocabulary (signature) $\Sigma$ consisting of: a set of constant symbols $\mathcal{C}$ (e.g., “Room1”,“Patient2”); a set of variable symbols $\mathcal{V}$ (e.g., position “xx”, distance“dd”); a set of function symbols $\mathcal{F}$ (e.g., $\text{neighborOf}(x)$”), each with an associated arity $n\in\mathbb{N}$ denoted by $f/n$ where $f\in\mathcal{F}$; and a set of predicate symbols $\mathcal{P}$ (e.g., $\text{Asleep}(x)$” or $\text{Nursing}(x,y)$”), also with associated arities $P/n$ where $P\in\mathcal{P}$. The set of terms $\mathcal{T}(\Sigma)$ over $\Sigma$ is then defined recursively: $\mathcal{C}\subset\mathcal{T}(\Sigma)$, $\mathcal{V}\subset\mathcal{T}(\Sigma)$; and if $f/n\in\mathcal{F}$ and $t_{1},\dots,t_{n}\in\mathcal{T}(\Sigma)$, then $f(t_{1},\dots,t_{n})\in\mathcal{T}(\Sigma)$ (e.g., “nurseOf(Patient1)”). The set of formulae $\mathcal{F}(\Sigma)$ over $\Sigma$ is then also defined recursively as follows. If $P/n\in\mathcal{P}$ and $t_{1},\dots,t_{n}\in\mathcal{T}(\Sigma)$, then $P(t_{1},\dots,t_{n})\in\mathcal{F}(\Sigma)$ is an atomic formula; if $\phi,\psi\in\mathcal{F}(\Sigma)$, then $\neg\phi,(\phi\land\psi),(\phi\lor\psi),(\phi\rightarrow\psi),(\phi\leftrightarrow\psi)\in\mathcal{F}(\Sigma)$, and if $\phi\in\mathcal{F}(\Sigma)$ and $x\in\mathcal{V}$, then $\forall x\phi$ and $\exists x\phi\in\mathcal{F}(\Sigma)$ (e.g., $\forall x(\text{Man}(x)\rightarrow\text{Mortal}(x))$”).
@@ 4.2.2 Planning
Conceptually, classical planning could be viewed as a special case of MPC where (i) the transition dynamics ($\mathcal{T}(s,a)$) are discrete and deterministic, (ii) the cost function prioritizes goal achievement (typically 0 if $x\models\mathcal{G}$ otherwise $\infty$), and (iii) the optimization horizon spans until the goal is reached.
@@ 4.4 Planning (and RL) with (Graph and) Relational Neural Networks
[…]
@@ 5.0.1 Fuzzy Sets
where $\mu_{\tilde{A}}(x)=1$ indicates full membership, $\mu_{\tilde{A}}(x)=0$ indicates non-membership, and $0<\mu_{\tilde{A}}(x)<1$ represents a partial membership. Common examples of membership functions include triangular, trapezoidal, Gaussian, and sigmoid functions bai2006fundamentals .
@@ 5.0.2 Fuzzy Logic
Fuzzy logic then extends classical Boolean logic by incorporating degrees of truth. In fuzzy logic, the truth value of a statement is again not restricted to binary values (0 or $1$) but can take any value in the interval $[0,1]$. To that aim, fuzzy logic operators extend classical logical operators to operate on the fuzzy sets bai2006fundamentals .
@@ 5.0.3 Fuzzy control
[…]
@@
$y^{*}=\frac{\int y\cdot\mu_{\tilde{B}_{\text{agg}}}(y)\,dy}{\int\mu_{\tilde{B}_{\text{agg}}}(y)\,dy}.$ (43)
@@
[…]
@@ 6 Integrating the Symbolic and Control Layers
[…]
Figure 3: High-level Integration Schema: A learning module (bottom) collects data from the robot to update parameters of the decision-making modules (top).
[…]
Figure 4: A temporal model of integrating the Planning and Control Layers. The slow-timescale Scheduler assigns tasks (e.g., Task A) which spawn Planning Problem(s). The discrete Planner issues actions $\hat{a}_{t}$ that trigger the fast-timescale MPC. Continuous physical states $\hat{z}$ map back to discrete symbolic states $\bar{s}$ via (de-)fuzzification ($\mu$). Possible discrepancies prompt real-time re-planning.
@@ 6.3 Action-Triggered MPC
$\quad h\left(z_{\tau(k)},v_{\tau(k)};p^{h}(a_{t})\right)\leq 0,\quad v_{\tau(k)}\in\tilde{A}(p^{a}(a_{t}))\,$$\quad$ (52c)
@@
[…]
Figure 5: Action-Triggered RLMPC Loop: The receding horizon principle. At each fast-timescale interval ($t,t+h,t+2h,\dots$), an optimal control problem is solved predicting the state trajectory over horizon $H$ (dashed lines). Only the first control action $\hat{v}(\tau)$ is applied, generating the true closed-loop trajectory (solid blue line). The RL Critic observes this realized performance and updates the cost parameters $\theta$.
@@
In addition, a real-time reward $\hat{R}^{e}(t+1)$ is observed externally at the conclusion of the problem, together with the evaluation of the action closed loop cost functional $\hat{R}^{o}(t+1)$. This quantity $\hat{R}^{e}(t+1)$ represents mission-relevant information that occurred in the course of the operation. For instance, if, in performing one of the control actions, the robot observed negative emotional affect in the human, this can enter into $\hat{R}^{e}(t+1)$ in a manner so as to inform the upper layer, and ultimately the task can be planned to be performed more agreeably to the human, that is, without action $\hat{a}_{t}$. This action-triggered receding horizon procedure, augmented by the RL critic updates to the cost parameters, is summarized in Figure 5.
@@ 6.4 Integrating State representations
where $p_{\text{RobotA}}$ and $p_{\text{LocationB}}$ are their respective positions, and $\sigma$ controls the smoothness of the transition. Note the truncation of the distribution: if the robot is within distance $\bar{d}^{0}_{At}$ of LocationB, we declare that it is At that location with certainty, and if it is a distance at least $\bar{d}^{1}_{At}$ away, then he is crisply declared to not be at the location, with a fuzzy transition function in the annulus defined by the second indicator in the expression. This translation mechanism is visualized in Figure 6.
Continuous Space($z_{t}\in\mathbb{R}^{2}$) Fuzzification($\mu_{L_{m}}$) Logic Atoms($s_{t}^{\mu}\in[0,1]^{M}$)
At(A) 0.9
At(B) 0.1
[…]
Figure 6: (De-)Fuzzification Layer: The translation from continuous physical coordinates (left), through fuzzy membership functions (middle), to discrete logic atoms (right) for the planner.
@@ 6.5 Updating the Planning State and Model
6.6 Updating the MPC Control Parametrization
In addition to, or instead of, updating the model of the state transition that operates at the planning layer, that is $p^{S}_{\psi}(\theta^{\psi};s_{t},a_{t})$ from the predicted actual state mismatch ($\bar{s}_{t+1}$ verus $\hat{s}_{t+1}$) and the intermediate cost $\Phi^{C}(\theta^{C};s_{a},a_{t})$ through observed rewards $\hat{C}_{t}$, we can also use the mismatch to update the action to MPC map $a_{t}\to(p^{w},p^{g},p^{f},p^{h},p^{a})$. That is, if the expected and desired final state is not reached upon triggering the MPC associated with action $a_{t}$, and/or there is unexpected reward or cost, then this information can be used to adjust the specific MPC problem generated by $a_{t}$.
We can compute the continuous form of the final state mismatch as:
$\hat{E}^{\psi}_{t}:=T^{\hat{s}_{t+1}}\bigwedge_{m}\mu_{L_{m}}(\hat{z}(t+1))-T^{\bar{s}_{t+1}}\bigwedge_{m}\mu_{L_{m}}(\hat{z}(t+1))$
that is, the fuzzy relaxation of $s_{t+1}$ as far as activation of the real state $\hat{s}_{t+1}$ compared to the originally predicted state $\bar{s}_{t+1}$. Through differentiation, we seek to obtain a correspondence between the discrete state transition through the fuzzy relaxation of the final state map given the final continuous state $\hat{z}(t+1)$ to the optimal MPC action that generated $\hat{z}_{t+1}$, through the parameters defining the MPC action as set by $a_{t}$.
There is also the reward mismatch between the actual realized reward through the low level state operation and the predicted cost associated with this action,
$\hat{E}^{C}_{t}:=\hat{R}_{t+1}-\bar{C}_{t+1}$
Now, applying the chain rule, we target the terminal cost in the MPC that we denoted by $W(z_{(t+1)};p^{w}(a_{t}))$ for adjustment as far as final state mismatch and the intermediate cost $G(z_{\tau(k)},v_{\tau(k)};p^{g}(a_{t}))$ for the reward mismatch, with learning rate $\alpha^{M}_{t}$ that may depend on the stage iteration $t$:
$\begin{array}[]{l}p^{w}(a_{t})\leftarrow p^{w}(a_{t})-\alpha^{M,w}_{t}\hat{E}^{\psi}_{t}\frac{d\hat{E}^{\psi}}{dp^{w}(a_{t})}\\ \frac{d\hat{E}^{\psi}}{dp^{w}(a_{t})}=\frac{d\hat{E}^{\psi}}{d\hat{z}_{t+1}}\frac{d\hat{z}_{t+1}}{dp^{w}(a_{t})}\\ p^{g}(a_{t})\leftarrow p^{g}(a_{t})-\alpha^{M,g}_{t}\hat{E}^{C}_{t}\sum\limits_{\tau=t+h}^{t+1-h}\left[\frac{dG}{d(z,v)}\frac{d(z_{\tau},v_{\tau})}{dp^{g}(a_{t})}+\frac{\partial G}{\partial p^{g}(a_{t})}\right]\end{array}$
MPC Sensitivities
We now describe how to compute the sensitivity of the fuzzy function $\mu_{L_{m}}(z^{\star}_{t+1})$ with respect to the MPC parameters $p_{t}$, where $z^{\star}_{t+1}$ is the end state of a simulation of the system in closed-loop with the MPC controller. These sensitivities will be useful to adjust the MPC parameters at the planning layer to ensure that the MPC scheme achieves the goals it is assigned.
We assume that the MPC scheme is run in a possibly complex but differentiable simulation model
$z^{\star}_{t+h(i+1)}=f^{\star}(z^{\star}_{t+hi},u_{t+hi}),\quad u_{t+hi}=\pi^{\mathrm{MPC}}(z^{\star}_{t+hi},p_{t})$ (60)
where $\pi^{\mathrm{MPC}}$ is the MPC policy, and $p_{t}$ are the MPC parameters selected by the supervisory layer, i.e.
$\pi^{\mathrm{MPC}}(z^{\star}_{t+hi},p_{t})=u_{0}^{\star}$ (61)
where $u_{0}^{\star}$ is the first element of the optimal input sequence $u^{\star}$ defined by solving the Parametric Nonlinear Program
$\min_{x,u}$ $\quad T(x_{N},p_{t})+\sum_{k=0}^{N-1}L(x_{k},u_{k},p_{t})$ (62a)
$\mathrm{s.t.}$ $\quad x_{k+1}=f^{\star}(x_{k},u_{k},p_{t}),\qquad x_{0}=z_{t+hi}^{\star}$ (62b)
$\quad h(x_{k},u_{k},p_{t})\geq 0$ (62c)
parametrized with the current simulation state $z^{\star}_{t+hi}$ and parameters $p_{t}$. We then establish how to differentiate a closed-loop simulation $\hat{z}^{\star}_{t,\ldots,t+h\tau}$ with respect to the MPC parameters $p_{t}$. Two approaches are useful, depending on the end-goal of the simulation differentiation: the forward and adjoint modes.
  • Forward mode This mode is useful if we are interested in the “overall” sensitivity of the simulation. In that mode, we make a forward pass on the simulation as:
    $\frac{\mathrm{d}z^{\star}_{t+h(i+1)}}{\mathrm{d}p_{t}}=\frac{\partial f^{\star}}{\partial z^{\star}_{t+hi}}\frac{\mathrm{d}z^{\star}_{t+hi}}{\mathrm{d}p_{t}}+\frac{\partial f}{\partial u_{t+hi}}\frac{\partial\pi^{\mathrm{MPC}}}{\partial p_{t}},\qquad\frac{\mathrm{d}z^{\star}_{t}}{\mathrm{d}p_{t}}=0$ (63)
    where all expressions are evaluated at $z^{\star}_{t+hi}$, $u_{t+hi}=\pi^{\mathrm{MPC}}(z^{\star}_{t+hi},p_{t})$ and $p_{t}$ alongside the forward simulations (60).
  • Adjoint mode If we are only interested in the sensitivity of the scalar quantity $\mu_{L_{m}}(z^{\star}_{t+h\tau})$ at the end of the simulation, then the adjoint-mode sensitivities is typically more effective. In that mode, the sensitivity is computed backward in time after the forward simulations (60) have been performed. We then define the adjoint variable:
    $\lambda^{i}=\frac{\mathrm{d}\mu_{L_{m}}(z^{\star}_{t+h\tau})}{\mathrm{d}z^{\star}_{t+hi}}^{\top},$ (64)
    whose dynamics read as
    $\lambda^{i-1}=\left.\frac{\partial f}{\partial z^{\star}_{t+(i-1)h}}^{\top}\right|_{z^{\star}_{t+hi},p_{t},u_{t+hi}}\cdot\lambda^{i},\qquad\lambda^{\tau}=\frac{\partial\mu_{L_{m}}(z^{\star}_{t+h\tau})}{\partial z^{\star}_{t+h\tau}}^{\top}.$ (65)
    Finally, the sensitivity of $\mu_{L_{m}}(z^{\star}_{t+h\tau})=\mu_{L_{m}}(z^{\star}_{t+1})$ to $p_{t}$ reads as
    $\frac{\mathrm{d}\mu_{L_{m}}(z^{\star}_{t+1})}{\mathrm{d}p_{t}}^{\top}=\sum_{i=0}^{\tau}\left.\left(\frac{\partial f}{\partial u_{t+hi}}\frac{\partial\pi^{\mathrm{MPC}}}{\partial p_{t}}\right)^{\top}\right|_{z^{\star}_{t+hi},p_{t},u_{t+hi}}\cdot\lambda^{i},$ (66)
    where $u_{t+hi}=\pi^{\mathrm{MPC}}(z^{\star}_{t+hi},p_{t})$.
Both the forward and adjoint-mode methods require computing the MPC policy sensitivity $\frac{\partial\pi^{\mathrm{MPC}}}{\partial p_{t}}$ at every simulation step $i$. It can be computed as follows.
Consider $w$ collecting the state-input variables $x,u$ of MPC (62), and
$w^{\star}(z^{\star}_{t+hi},p_{t})$ the optimal solution to (62) for a give initial condition $z^{\star}_{t+hi}$ and MPC parameters $p_{t}$. Consider in addition function $G(w,z^{\star}_{t+hi},p_{t})$ collecting the equality constraints (62b) and function $H(w,z^{\star}_{t+hi},p_{t})$ collecting the inequality constraints (62c), and the Lagrange function:
$\mathcal{L}(\zeta,z^{\star}_{t+hi},p_{t})=$ $T(x_{N},p_{t})+\sum_{k=0}^{N-1}L(x_{k},u_{k},p_{t})+\lambda^{\top}G(w,z^{\star}_{t+hi},p_{t})$ (67)
$+\mu^{\top}H(w,z^{\star}_{t+hi},p_{t})$
where $\zeta=\left\{w,\lambda,\mu\right\}$ collects the primal-dual variables associated to (62), and $\lambda,\mu$ are the dual variables associated to the equality and inequality constraints (62b)-(62c). The primal-dual Interior-Point KKT conditions underlying a solution to (62) read as:
$\kappa_{\nu}(\zeta,z^{\star}_{t+hi},p_{t})=\left[\begin{array}[]{c}\nabla_{w}\mathcal{L}(\zeta,z^{\star}_{t+hi},p_{t})\\ G(w,z^{\star}_{t+hi},p_{t})\\ \mu\circ H(w,z^{\star}_{t+hi},p_{t})+\nu\end{array}\right]=0$ (68d)
where $\circ$ is the element-wise multiplication, and where $\nu\geq 0$ is the barrier parameter underlying primal-dual Interior Point methods. The NLP is typically solved with $\nu$ small. Assuming that the objective (62a) is at least twice continuously differentiable, and that the constraints (62b)-(62c) are at least continuously differentiable, the Implicit Function Theorem (IFT) guarantees the existence of the solution sensitivities $\frac{\partial\zeta^{\star}}{\partial p_{t}}$ at $p_{t}$ if the linear independence constraint qualification, second-order sufficient conditions, and strict complementarity are satisfied at the solution $\zeta^{\star}$. Under these conditions, the IFT furthermore implies that the solution sensitivity can be computed by solving the linear system:
$\left.\frac{\partial\kappa_{\nu}(\zeta,z^{\star}_{t+hi},p_{t})}{\partial\zeta}\frac{\mathrm{d}\zeta^{\star}}{\mathrm{d}p_{t}}+\frac{\partial\kappa_{\nu}(\zeta,z^{\star}_{t+hi},p_{t})}{\partial p_{t}}\right|_{\zeta=\zeta^{\star}}=0$ (69)
for $\frac{\mathrm{d}\zeta^{\star}}{\mathrm{d}p_{t}}$. The sensitivity of the MPC prediction can then be extracted from that solution. Note that the linear system (69) is well-posed if the MPC scheme achieves the LICQ and SOSC conditions.
Recent software developments implement algorithms to both solve the MPC problem and to compute the sensitivity of its solution efficiently, see e.g. frey2025differentiable . The algorithmic parameter $\tau$ can be used to provide a solution manifold to the NLP that is smooth everywhere, such that the sensitivities are well defined for all $p_{t}$. Note that the approximation introduced by $\nu>0$ is of the order $\nu$.
@@ 6.7 Fuzzy control relaxation alternative
[…]